OpenStudy (anonymous):

A sequence is defined recursively using the equation f(n + 1) = f(n) – 8. If f(1) = 100, what is f(6)?

OpenStudy (freckles):

try to find f(2) first then f(3) then f(4) then f(5) then finally f(6)

OpenStudy (freckles):

find f(2) by replacing the n's with 1 f(n+1)=f(n)-8 f(1+1)=f(1)-8 f(2)=f(1)-8=?

OpenStudy (anonymous):

i am not sure how to do that

OpenStudy (freckles):

f(1) is given...

OpenStudy (freckles):

do you see f(1)=100 above?

OpenStudy (freckles):

replace f(1) with 100 so f(2)=100-8=?

OpenStudy (anonymous):

oh that make sense now thanks

OpenStudy (freckles):

now try finding f(3) by replacing the n's with 2 instead

OpenStudy (anonymous):

f(3)=182

OpenStudy (freckles):

\[f(n+1)=f(n)-8 \\ \\ \text{ replace } n \text{ with 2 } \\ f(2+1)=f(2)-8 \\ f(3)=f(2)-8 \\ \text{ remember we found } f(2) \text{ to be 92 } \\ f(3)=92-8=?\]

OpenStudy (anonymous):

84

OpenStudy (freckles):

you should notice that each time we do these outputs are decreasing by 8 units we started with f(1)=100 then we got f(2)=92 then we got f(3)=84 so we should get f(4)=?

OpenStudy (freckles):

just keep following the pattern take 8 away from previous output

OpenStudy (anonymous):

ok got it

OpenStudy (freckles):

and f(6) isn't too far away :p

OpenStudy (anonymous):

cool I got f(6)=60